OpenStudy (anonymous):

find the coordinates of the vertices and foci and the equations of the asymptotes for the hyperbola: x^2/9-y^2/4=1

OpenStudy (anonymous):

For any hyperbola, x^2/a^2 - y^2/b^2=1 the coordinate of vertices would be ( a,0) and (-a,0), focus would be (ae,0) and (-ae,0) where e= \[\sqrt[2]{1+b^{2}/a ^{2}}\] and the asymptotes are y= bx/a and y = -bx/a. IN your question, a=3 and b=4, good luck